# Stellar-gas kinematic misalignments in eagle: lifetimes and longevity of misaligned galaxies

### v1

@@ 6.1 Misalignment distributions
Figure 4: Evolution of two galaxies that experience long relaxation periods. Left: Galaxy A displaying an unstable misalignment matching our criteria with $t_{\rm{relax}}\approx 3.78$ Gyr $\approx 15.0\bar{t}_{\rm{dyn}}\approx 3.5\bar{t}_{\rm{torque}}$. Right: Galaxy B shows an initial unstable misalignment and relaxation into the stable counter-rotating regime with $t_{\rm{relax}}\approx 0.35$ Gyr $\approx 2.0\bar{t}_{\rm{dyn}}\approx 1.1\bar{t}_{\rm{torque}}$, before the cumulative effects of star formation induce a second unstable misalignment driven by a stellar angular momentum flip back into a co-rotating state with $t_{\rm{relax}}\approx 2.55$ Gyr $\approx 12.0\bar{t}_{\rm{dyn}}\approx 7.5\bar{t}_{\rm{torque}}$. In order, the rows show: (1) the 3D misalignment angle between stars and gasSF (black) within $r_{50}$ ($1\sigma$ error bars shown in grey) and the 3D misalignment angle between the stars/gasSF at $r_{50}$ and the DM at 30 pkpc, with the stable regime ($<20^{\circ}$ and $>160^{\circ}$) and aligned regime ($<30^{\circ}$ and $>150^{\circ}$) highlighted by the horizontal grey shaded regions, (2) the dynamical time (solid green line) and torquing time (solid black line) of the system, (3) the total gas inflow (solid green line) and outflow (dashed red line) rate within $r_{50}$ and the SFR multiplied by a factor of 10 (dashed orange line) within $2r_{50}$, (4) the stellar mass (solid red line) and gasSF mass (dotted blue line) within $2r_{50}$, (5) the stellar co-rotational energy (red solid line) that we use as a proxy for morphology (with the horizontal black line marking the boundary between ETGs and LTGs), the gasSF co-rotational energy within $r_{50}$ (blue dash-dot line), and the stellar ellipticity (purple dashed line). Mergers are indicated by the vertical grey dashed line and occur at $t_{\rm{lookback}}=3.61$ ($\mu_{\rm{*}}=0.32$) for Galaxy A, and $t_{\rm{lookback}}=6.88$ ($\mu_{\rm{*}}=0.30$) for Galaxy B.
@@ 6.2 Formation and relaxation process of stellar-gas kinematic misalignments
To highlight an example of a long-lived unstable misalignment and to illustrate our method involving the identification and extraction of unstable misalignment timescales, we provide examples of two galaxy evolutions in Figure 4. Both galaxies experience at least one unstable misalignment with a relaxation time $t_{\rm{relax}}>2$ Gyr. These galaxies were selected for their interesting stellar-gasSF kinematic evolution and the longevity of their unstable misalignments (as detailed below). We stress that the exceptionally long relaxation times of the two selected galaxies tend to be a rarity among our sample, as we discuss in detail in Section 6.3.
@@
Galaxy A (left panels of Figure 4) is a relatively low-mass galaxy with $M_{*}\approx 8\times 10^{9}$ M at $z=0$ and with a single long-duration unstable misalignment starting at a lookback time $t_{\mathrm{lookback}}=4.38$ Gyr. Around 1 Gyr before the unstable misalignment, this galaxy experiences a close encounter with a lower-mass satellite which decouples the angular momentum of the galaxy from its wider DM halo (as seen from the DM-stars and DM-gasSF misalignment angles illustrated in the top-left panel of Figure 4). The galaxy remains marginally misaligned with its dark matter halo for the remainder of its evolution, with an average stellar-DM misalignment angle of $\approx 30^{\circ}$. A second close fly-by strips a large amount of gas from the satellite which accretes onto the central galaxy (seen as a sharp increase in gas inflow rate), forming the unstable misalignment. This unstable misalignment remains at $\psi_{\rm{3D}}\approx 60^{\circ}$ before beginning to show signs of relaxation at $t_{\mathrm{lookback}}\approx 2.5$ Gyr, eventually settling into the stable co-rotating regime at $t_{\mathrm{lookback}}=0.60$ Gyr with a $t_{\rm{relax}}\approx 3.78$ Gyr. Expressed in terms of time-averaged $\bar{t}_{\rm{dyn}}$ and $\bar{t}_{\rm{torque}}$ we find $t_{\rm{relax}}\approx 15.0\bar{t}_{\rm{dyn}}\approx 3.5\bar{t}_{\rm{torque}}$, indicating this unstable misalignment has persisted for significantly longer than predicted by a purely dynamical settling time.
@@
While in the process of relaxing, the galaxy undergoes a major merger with the previously gas-stripped satellite at $t_{\mathrm{lookback}}=3.61$ Gyr (horizontal dashed line in the left panels of Figure 4) with a stellar mass ratio of $\mu_{\rm{*}}=0.32$. While this merger does not substantially change $t_{\rm{dyn}}$, it causes a significant morphological transformation from an LTG to an ETG (seen as a decrease in $\kappa_{\rm{co}}^{*}$; bottom-left panel of Figure 4) and a decrease in the stellar ellipticity. This drives a significant and permanent increase in $t_{\rm{torque}}$ and helps drive an overall longer relaxation time. This behaviour is not uncommon among other long-duration unstable misalignments. We find a $>20\%$ decrease in stellar ellipticity between pre- and post-unstable misalignment formation is seen in $\approx 33\%$ of $t_{\rm{relax}}>2$ Gyr relaxations. Despite the increase in $\bar{t}_{\rm{torque}}$ to $1-2$ Gyr, the galaxy remains unable to fully relax for another $\approx 3$ Gyr.
@@
Galaxy B (right panels of Figure 4) shows a very different overall kinematic evolution. Analogous to Galaxy A, this galaxy becomes misaligned with respect to its DM halo shortly before a gas-rich major merger ($\mu_{\rm{*}}=0.30$) at $t_{\mathrm{lookback}}=6.88$ Gyr. This merger feeds an inflow of gas to the inner $r_{50}$, disrupting the existing gasSF and causing it to settle quickly into a retrograde disc in the short timescale of $t_{\rm{relax}}\approx 0.35$ Gyr $\approx 2.0\bar{t}_{\rm{dyn}}\approx 1.1\bar{t}_{\rm{torque}}$. The galaxy remains in a counter-rotating configuration for $\sim 2$ Gyr. During this time, the retrograde cold gas forms a new counter-rotating component of younger stars alongside the existing older co-rotating stars. This is seen from the steady decrease in gasSF mass and a corresponding increase in stellar mass, together with an increase in stellar co-rotational energy transforming the system back into a rotationally-supported LTG over $\approx 4$ Gyr. Starting at $t_{\mathrm{lookback}}\approx 4.4$ Gyr, we begin to see an increase in the $1\sigma$ uncertainty in the stellar-gasSF misalignment angle. This marks the turning point at which the angular momentum of the two stellar components compete over dominance, beginning the second unstable misalignment and realignment into a co-rotational state. This stellar net-angular momentum flip (traced both by the stellar-gasSF angle and stellar-DM angle; see the top-right panel of Figure 4) occurs gradually, as reflected by the long relaxation time $t_{\rm{relax}}\approx 2.55$ Gyr $\approx 12.0\bar{t}_{\rm{dyn}}\approx 7.5\bar{t}_{\rm{torque}}$, starting at $t_{\mathrm{lookback}}=4.38$ Gyr.
Figure 5: Stellar and gasSF mass distributions within $2r_{50}$ of our $z=0.1$ unstable misalignment distribution sample (orange/light blue shaded regions) and our relaxation sample (red/blue lines) between $10<z<$ 1. We find our relaxation sample is broadly representative of the sample at $z=0.1$ in terms of stellar mass, with a slight bias toward higher gasSF masses.
@@ 6.3 Relaxation timescale distributions
Using our selection criteria (see Section 5.3.1) we extract a sample of 3154 unstable misalignment relaxations between $10<z<$ 1. We include each instance of an unstable misalignment separately, even if these occur in the same galaxy (such as Galaxy B described in Section 6.2). Median values and statistics are found in Table 2 and Table 3.
@@ 6.3.1 The total galaxy population
Galaxies in eagle display a variety of relaxation times and pathways (see Figure 7 below). Figure 6 shows the distributions of relaxation times (top), and normalised relaxation times in terms of $t_{\rm{relax}}/\bar{t}_{\rm{tdyn}}$ (middle), and as $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (bottom) for our total sample of unstable misalignment relaxations. The majority of unstable misalignments in our sample tend to be fairly short-lived, with median relaxation times of $t_{\rm{relax}}=0.50$ Gyr $=3.2\bar{t}_{\rm{dyn}}=1.4\bar{t}_{\rm{torque}}$.
@@
Figure 6: Distributions of relaxation times of our total sample of 3154 unstable misalignment relaxations for $10<z<$ 1. Relaxations are coloured according to the relaxation path: co-rotating to co-rotating (blue), co-rotating to counter-rotating (green), counter-rotating to counter-rotating (orange), and counter-rotating to co-rotating (red). Inset graphs show the same total distribution on a log percentage axis, with a line of best fit (dashed purple). Errors are given as Poisson uncertainties. We find a broadly log-linear distribution (inset graphs) of relaxation time distributions, with the majority of unstable misalignments relaxing within shorter timescales but with a small population of longer-lived misalignments.
Focusing on the tail-end of the distributions, we find a moderate population ($\approx 20.1\%$) of unstable misalignments persisting for $t_{\rm{relax}}>1$ Gyr, with only a handful of examples exceeding $\gtrsim 20\bar{t}_{\rm{dyn}}$. Additionally, we find $159$ ($5.0\%$) instances of unstable misalignments with $t_{\rm{relax}}>2$ Gyr (two of which are shown and discussed in Section 6.2) that are approximately of similar duration to the galaxy investigated by van de Voort et al. (2015) in the fire simulations (Hopkins et al., 2014). The population of long-duration unstable misalignments becomes a little clearer when expressed as $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, with $\approx 19.9\%$ of relaxations lasting more than three times longer ($>3\bar{t}_{\rm{torque}}$) than expected from a purely dynamical relaxation (see Section 2). We note that while our relaxation selection method identifies two unstable misalignments with exceptionally enhanced relaxation timescales of $\gtrsim 25t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, closer inspection reveals these to be systems that spent considerable time ($>1.5$ Gyr) near the stable counter-rotating regime without settling. We attribute these outliers to the consequences of choosing a non-zero latency time.
@@
All $-0.71t_{\rm{relax}}-0.42$ $-0.09\bar{t}_{\rm{dyn}}-0.37$ $-0.20\bar{t}_{\rm{torque}}-0.31$
LTG $\rightarrow$ LTG $-1.16t_{\rm{relax}}-0.15$ $-0.10\bar{t}_{\rm{dyn}}-0.33$ $-0.18\bar{t}_{\rm{torque}}-0.36$
ETG $\rightarrow$ ETG $-0.66t_{\rm{relax}}-0.45$ $-0.08\bar{t}_{\rm{dyn}}-0.38$ $-0.21\bar{t}_{\rm{torque}}-0.30$

### v2

@@ 6.1 Misalignment distributions
Figure 4: Evolution of two galaxies that experience long relaxation periods. Left: Galaxy A displaying an unstable misalignment matching our criteria with $t_{\rm{relax}}\approx 3.78$ Gyr $\approx 15.0\bar{t}_{\rm{dyn}}\approx 3.5\bar{t}_{\rm{torque}}$. Right: Galaxy B shows an initial unstable misalignment and relaxation into the stable counter-rotating regime with $t_{\rm{relax}}\approx 0.35$ Gyr $\approx 2.0\bar{t}_{\rm{dyn}}\approx 1.1\bar{t}_{\rm{torque}}$, before the cumulative effects of star formation induce a second unstable misalignment driven by a stellar angular momentum flip back into a co-rotating state with $t_{\rm{relax}}\approx 2.55$ Gyr $\approx 12.0\bar{t}_{\rm{dyn}}\approx 7.5\bar{t}_{\rm{torque}}$. In order, the rows show: (1) the 3D misalignment angle between stars and gasSF (black) within $r_{50}$ ($1\sigma$ error bars shown in grey) and the 3D misalignment angle between the stars/gasSF at $r_{50}$ and the DM at 30 pkpc, with the horizontal dark and light grey shaded regions representing the stable regime ($20^{\circ}$ and $160^{\circ}$) and aligned/counter-rotating regime ($30^{\circ}$ and $150^{\circ}$), respectively
, (2) the dynamical time (solid green line) and torquing time (solid black line) of the system, (3) the total gas inflow rate at $r_{50}$ (solid green line) and the SFR multiplied by a factor of 10 (dashed orange line) within $2r_{50}$, (4) the stellar mass (solid red line) and gasSF mass (dotted blue line) within $2r_{50}$, (5) the stellar co-rotational energy (red solid line) that we use as a proxy for morphology (with the horizontal black line marking the boundary between ETGs and LTGs), the gasSF co-rotational energy within $r_{50}$ (blue dash-dot line), and the stellar ellipticity (purple dashed line). Mergers are indicated by the vertical grey dashed line and occur at $t_{\rm{lookback}}=3.61$ ($\mu_{\rm{*}}=0.32$) for Galaxy A, and $t_{\rm{lookback}}=6.88$ ($\mu_{\rm{*}}=0.30$) for Galaxy B.
The longevity of counter-rotating systems in eagle is also in agreement with results obtained using other cosmological hydrodynamical simulations (e.g. Starkenburg et al., 2019; Cenci et al., 2024). Starkenburg et al. (2019) use the IllustrisTNG (Nelson et al., 2019) hydrodynamical simulations to investigate the longevity of counter-rotating systems, finding these to be long-lasting relative to other orientations. Similarly, Cenci et al. (2024) investigated the formation of a long-lived counter-rotating galaxy within the FIREbox simulations (Feldmann et al., 2023). Following a starburst induced by the interaction with a nearby low-mass satellite and the subsequent rapid depletion of co-rotating gas, a long-lived ($\sim 4$ Gyr) counter-rotating gas disc is able to form from the accretion of stripped misaligned gas. We note that carrying out a detailed analysis of this behaviour in eagle is beyond the scope of this work.
@@ 6.2 Formation and relaxation process of stellar-gas kinematic misalignments
Figure 5: Illustration of the evolution of two galaxies shown in Figure 4 in five key stages. The net angular momentum vectors of the stellar (yellow) and gasSF (blue) are represented by arrows. Top: evolution of Galaxy A beginning with (I) an initial flyby of a gas-rich satellite that causing a stellar-DM misalignment; (II) the formation of an unstable misalignment from the accretion of gas stripped from the satellite during a second fly-by; (III) merging of the satellite galaxy with the central galaxy and the resulting morphological transformation of the galaxy to a more spheroidal stellar mass distribution, while maintaining the unstable misalignment through continued accretion of stripped gas; (IV) a decrease in the inflow rate allows the gas disc to begin precessing and relaxing into the galactic plane; (V) the gas disc becomes aligned with the stellar angular momentum. Bottom: evolution of Galaxy B beginning with (I) a gas-rich merger that disrupts the existing gas disc; (II) the quick formation of a large misaligned gas disc that quickly settles to become counter-rotating; (III) a new generation of counter-rotating stars begin to form from the counter-rotating gas which begins to build a new counter-rotating stellar disc; (IV) the stellar angular momentum eventually flips, manifesting as a ‘pseudo-relaxation’; (V) the counter-rotating stars dominate the net angular momentum and the galaxy appears as aligned.
@@
To highlight an example of a long-lived unstable misalignment and to illustrate our method involving the identification and extraction of unstable misalignment timescales, we provide examples of two galaxy evolutions in Figure 4. Illustrations of these evolutions are provided in Figure 5 showing five key stages in each galaxy’s evolution (labelled I-V). We will refer to these stages throughout this subsection. Both galaxies experience at least one unstable misalignment with a relaxation time $t_{\rm{relax}}>2$ Gyr. These galaxies were selected for their interesting stellar-gasSF kinematic evolution and the longevity of their unstable misalignments (as detailed below). We stress that the exceptionally long relaxation times of the two selected galaxies tend to be a rarity among our sample, as we discuss in detail in Section 6.3.
@@
Galaxy A (left panels of Figure 4) is a relatively low-mass galaxy with $M_{*}\approx 8\times 10^{9}$ M at $z=0$ and with a single long-duration unstable misalignment starting at a lookback time $t_{\mathrm{lookback}}=4.38$ Gyr. (I) Around 1 Gyr before the unstable misalignment, this galaxy experiences a close encounter with a lower-mass ($M_{*}\approx 3\times 10^{9}$ M) satellite which decouples the angular momentum of the galaxy from its wider DM halo (see also e.g. Velliscig et al., 2015). This is seen as a steep increase in the DM-stars and DM-gasSF misalignment angle illustrated in the top-left panel of Figure 4. The galaxy remains marginally misaligned with its dark matter halo for the remainder of its evolution, with an average stellar-DM misalignment angle of $\approx 30^{\circ}$. (II) A second close fly-by strips a large amount of gas from the satellite. Most of this gas accretes onto the central galaxy (seen as a sharp increase in gas inflow rate), disrupting the existing co-rotating gas disc (seen as a decrease in $\kappa_{\rm{co}}^{\rm{SF}}$), and quickly forms an unstable misalignment. (III) Ongoing accretion of stripped gas immediately after the misalignment formation is able to sustain the angular momentum of the misaligned gas disc, maintaining a steady misalignment angle of $\psi_{\rm{3D}}\approx 60^{\circ}$ until $t_{\mathrm{lookback}}\sim 2.5$ Gyr. (IV) This is followed by a decrease in the gas inflow rate and the onset of differential precession about the stellar angular momentum axis. The unstable misalignment begins to relax toward the galactic plane, (V) eventually settling into the stable co-rotating regime at $t_{\mathrm{lookback}}=0.60$ Gyr with a $t_{\rm{relax}}\approx 3.78$ Gyr. Expressed in terms of time-averaged $\bar{t}_{\rm{dyn}}$ and $\bar{t}_{\rm{torque}}$ we find $t_{\rm{relax}}\approx 15.0\bar{t}_{\rm{dyn}}\approx 3.5\bar{t}_{\rm{torque}}$, indicating this unstable misalignment has persisted for significantly longer than predicted by a purely dynamical settling time. We note that the behaviour of this misalignment post-formation is similar to that studied by van de Voort et al. (2015) using the fire (Hopkins et al., 2014) cosmological zoom-in simulations.
@@
While in the process of relaxing, the galaxy undergoes a major merger with the previously gas-stripped satellite at $t_{\mathrm{lookback}}=3.61$ Gyr (horizontal dashed line in the left panels of Figure 4) with a stellar mass ratio of $\mu_{\rm{*}}=0.32$. While this merger does not substantially change $t_{\rm{dyn}}$, it causes a significant morphological transformation from an LTG to an ETG (seen as a decrease in $\kappa_{\rm{co}}^{*}$; bottom-left panel of Figure 4) and a decrease in the stellar ellipticity (see panel III in Figure 5). This drives a significant and permanent increase in $t_{\rm{torque}}$ and helps drive an overall longer relaxation time. This behaviour is not uncommon among other long-duration unstable misalignments. We find a >20>20 per cent decrease in stellar ellipticity between pre- and post-unstable misalignment formation is seen in $\approx 33$ per cent of $t_{\rm{relax}}>2$ Gyr relaxations. Despite the increase in $\bar{t}_{\rm{torque}}$ to 121-2 Gyr, the galaxy remains unable to fully relax for another $\approx 3$ Gyr. This is due to a small gas accretion event at $t_{\mathrm{lookback}}\sim 1.5$ Gyr, although the accretion rate is insufficient to prevent further differential precession and eventual relaxation into stable co-rotation.
Galaxy B (right panels of Figure 4) shows a very different overall kinematic evolution. (I) Analogous to Galaxy A, this galaxy becomes misaligned with respect to its DM halo shortly before a gas-rich major merger ($\mu_{\rm{*}}=0.30$) at $t_{\mathrm{lookback}}=6.88$ Gyr. (II) This merger feeds an inflow of gas to the inner $r_{50}$, disrupting the existing gasSF and causing it to settle quickly into a retrograde disc in the short timescale of $t_{\rm{relax}}\approx 0.35$ Gyr $\approx 2.0\bar{t}_{\rm{dyn}}\approx 1.1\bar{t}_{\rm{torque}}$. (III) The galaxy remains in a counter-rotating configuration for $\sim 2$ Gyr. During this time, the retrograde cold gas forms a new counter-rotating component of younger stars alongside the existing older co-rotating stars. This is seen from the steady decrease in gasSF mass and a corresponding increase in stellar mass, together with an increase in stellar co-rotational energy transforming the system back into a rotationally-supported LTG over $\approx 4$ Gyr. (IV) Starting at $t_{\mathrm{lookback}}\approx 4.4$ Gyr, we begin to see an increase in the $1\sigma$ uncertainty in the stellar-gasSF misalignment angle. This marks the turning point at which the angular momentum of the two stellar components compete over dominance, beginning the second unstable misalignment and realignment into a co-rotational state. As the apparent relaxation was driven by a change in the stellar net-angular momentum rather than an explicit relaxation of the gas disc, this event can be considered a ’pseudo-relaxation’. This stellar net-angular momentum flip (traced both by the stellar-gasSF angle and stellar-DM angle; see the top-right panel of Figure 4) occurs gradually, as reflected by the long relaxation time $t_{\rm{relax}}\approx 2.55$ Gyr $\approx 12.0\bar{t}_{\rm{dyn}}\approx 7.5\bar{t}_{\rm{torque}}$, starting at $t_{\mathrm{lookback}}=4.38$ Gyr. (V) By $t_{\mathrm{lookback}}\approx 2.4$ Gyr the galaxy shows a clear alignment between the (now discy) stellar and gasSF components.
@@
Figure 6: Stellar and gasSF mass distributions within $2r_{50}$ of our $z=0.1$ unstable misalignment distribution sample (orange/light blue shaded regions) and our relaxation sample (red/blue lines) between $0<z<1$. We find our relaxation sample is broadly representative of the sample at $z=0.1$ in terms of stellar mass, with a slight bias toward higher gasSF masses.
@@ 6.3 Relaxation timescale distributions
Using our selection criteria (see Section 5.3.1) we extract a sample of 3154 unstable misalignment relaxations between $0<z<1$. We include each instance of an unstable misalignment separately, even if these occur in the same galaxy (such as Galaxy B described in Section 6.2). Median values and statistics are found in Table 2 and Table 3.
@@ 6.3.1 The total galaxy population
Galaxies in eagle display a variety of relaxation times and pathways. Figure 7 shows the distributions of relaxation times (top), and normalised relaxation times in terms of $t_{\rm{relax}}/\bar{t}_{\rm{tdyn}}$ (middle), and as $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$ (bottom) for our total sample of unstable misalignment relaxations. The majority of unstable misalignments in our sample tend to be fairly short-lived, with median relaxation times of $t_{\rm{relax}}=0.50$ Gyr $3.2\bar{t}_{\rm{dyn}}=1.4\bar{t}_{\rm{torque}}$.
@@
Figure 7: Distributions of relaxation times of our total sample of 3154 unstable misalignment relaxations for $0<z<1$. Relaxations are coloured according to the relaxation path: co-rotating to co-rotating (blue), co-rotating to counter-rotating (green), counter-rotating to counter-rotating (orange), and counter-rotating to co-rotating (red). Inset graphs show the same total distribution on a log percentage axis, with a line of best fit (dashed purple). Errors are given as Poisson uncertainties. We find a broadly log-linear distribution (inset graphs) of relaxation time distributions, with the majority of unstable misalignments relaxing within shorter timescales but with a small population of longer-lived misalignments.
Focusing on the tail-end of the distributions, we find a moderate population ($\approx 20.1$ per cent) of unstable misalignments persisting for $t_{\rm{relax}}>1$ Gyr, with only a handful of examples exceeding $\gtrsim 20\bar{t}_{\rm{dyn}}$. Additionally, we find $159$ ($5.0$ per cent) instances of unstable misalignments with $t_{\rm{relax}}>2$ Gyr (two of which are shown and discussed in Section 6.2) that are approximately of similar duration to the galaxy investigated by van de Voort et al. (2015) in the fire simulations (Hopkins et al., 2014). The population of long-duration unstable misalignments becomes a little clearer when expressed as $t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, with $\approx 19.9$ per cent of relaxations lasting more than three times longer ($3\bar{t}_{\rm{torque}}$) than expected from a purely dynamical relaxation (see Section 2). We note that while our relaxation selection method identifies two unstable misalignments with exceptionally enhanced relaxation timescales of $\gtrsim 25t_{\rm{relax}}/\bar{t}_{\rm{torque}}$, closer inspection reveals these to be systems that spent considerable time (>1.5>1.5 Gyr) near the stable counter-rotating regime without settling. We attribute these outliers to the consequences of choosing a non-zero latency time.
@@
All $71t_{\rm{relax}}-0.42 0.0$$09\bar{t}_{\rm{dyn}}-0.37$ $0.20\bar{t}_{\rm{torque}}-0.31$
LTG $\rightarrow$ LTG $1.16t_{\rm{relax}}-0.15$ $0.10\bar{t}_{\rm{dyn}}-0.33$ $0.18\bar{t}_{\rm{torque}}-0.36$
ETG $\rightarrow$ ETG $66t_{\rm{relax}}-0.45 0.0$$08\bar{t}_{\rm{dyn}}-0.38$ $0.21\bar{t}_{\rm{torque}}-0.30$